# Spectral Masking and Interpolation Attack (SMIA): A Black-box Adversarial Attack against Voice Authentication and Anti-Spoofing Systems

### v1

@@ 2.6.2. Black-box Attacks
In the domain of black-box adversarial attacks, Hai et al. proposed SiFDetectCracker, a framework targeting fake voice detection systems (CMs). Their method generates perturbations for speaker-irrelevant features (SiFs), specifically the background noise and the mute parts of an audio sample. The framework was evaluated against several advanced detectors, including RawNet2(Tak et al., 2021b), RawGAT-ST(Tak et al., 2021a), RawPC-Darts (Ge et al., 2021), and Deep4SNet (Ballesteros et al., 2021), using the ASVspoof 2019 dataset (et al., 2019). The attack proved highly effective, achieving an average success rate of 82.2% by showing that many detection systems are sensitive to these non-speech features. While the paper presents a novel attack vector, ablation studies showed that the attack’s performance drops significantly when either time or noise perturbations are removed, indicating a reliance on both factors. Furthermore, the transferability between models with different architectures was low, a limitation our work aims to address.
While the work by Kassis and Hengartner (Kassis and Hengartner, 2023) presents a novel methodology for bypassing voice authentication (VA) systems, its real-world applicability is limited by key limitations in its evaluation. Their results against combination systems rely on a cumulative metric, which is not descriptive enough to show the state or the actual effectiveness of their attack against these more complex defenses. Furthermore, the attack’s performance dropped significantly when tested against SOTA anti-spoofing systems, like RawPC-Darts (Ge et al., 2021) and RawGAT-ST (Tak et al., 2021a), raising doubts about its effectiveness against modern countermeasures. Finally, the designated-app attack vector relies on the significant assumption that an attacker can use a rooted phone to inject audio, a prerequisite that is often blocked by the security measures of financial applications.
Paper Year Tasks Attack Type System Target
Deepfake Voice Attacks
Saxena et al. (Shonesy, 2015) 2015 SV Black-box (Voice morphing) SV, Human listeners
Alali et al. (Alali and Theodorakopoulos, 2025) 2025 SV Black-box (Zero-shot TTS / VC) SV, Human listeners
Adversarial Spoofing Attacks
Liu et al. (Liu et al., 2019) 2019 CM White-box (FGSM/PGD) ASVspoof CM
Gomez-Alanis et al. (Gomez-Alanis et al., 2022a) 2022 SV + CM White-box (ABTN) ASVspoof systems
Gomez-Alanis et al. (Gomez-Alanis et al., 2022b) 2022 SV + CM White-box (GANBA) ASVspoof systems
Hai et al. (Hai et al., 2023) 2023 CM Black-box CM
Kassis et al. (Kassis and Hengartner, 2023) 2023 SV + CM Black-box (OTA spoofing) VA systems (SV + CM)
2.7. Threat Model
Our threat model considers an attacker whose primary goal is to gain unauthorized access to a victim’s account on a Speaker Verification (SV) system by mimicking their voice. The success of the attack is defined by a single successful authentication event.
The attacker operates under a black-box assumption, meaning they have no internal knowledge of the system’s architecture, including the specific Voice Authentication System (VAS) or Countermeasure (CM) models employed. The attacker’s capabilities are as follows:
  • They can collect voice samples of the victim from publicly available sources, requiring a minimum of 10 seconds of audio to initiate the attack.
  • They can generate a spoofed audio sample using a voice-cloning tool (e.g., Fish Speech APIs (Liao et al., 2024), methods used in ASVspoof 2019 (et al., 2019), or any similar tools).
  • The attacker can repeatedly submit the generated audio to the system’s authentication endpoint and observe the binary pass/fail outcome. This feedback is used to iteratively optimize the malicious audio sample to bypass both the VAS and CM defenses.
The attacker’s actions are considered stealthy, as a successful login does not interfere with or alert the legitimate user. The attack is applicable to both text-dependent and text-independent SV systems, though text-dependent systems present a higher barrier, requiring the attacker to also know the specific pass-phrase for the authentication session.
@@ 3. Methodology
A combined voice authentication and anti-spoofing system processes a given voice sample $x\in\mathbb{R}^{m}$ to perform two key tasks: speaker verification (SV) and spoofing detection (CM). The system grants access only when the sample is verified as belonging to the claimed identity and determined to be a genuine, live utterance.
@@
(1) $S(x,e)=\begin{cases}\text{accept}&\text{if }g(x,e)=\text{accept}\land f(x)=\text{real}\\ \text{reject}&\text{otherwise}\end{cases}$
@@ 3.1. Overview
As shown in Figure 2, the attack pipeline starts by generating a base spoofed audio. This is accomplished by feeding a victim’s voice sample and a target prompt into a voice synthesis model. Our framework is tool-independent—you can employ Fish Speech APIs (Liao et al., 2024), ASVSpoof challenge tools (et al., 2019), or any modern voice cloning system.
Second, this synthesized audio serves as the input to our core attack mechanism: the Spectral Masking and Iterative Attack (SMIA) system. SMIA operates within an iterative loop where it strategically modifies the audio using our proposed Spectral Masking and Interpolation technique. The resulting adversarial audio is submitted to the target system (VAS or CM), and the prediction outcome is used as feedback to update the SMIA parameters. This process is repeated, progressively refining the audio until it successfully deceives the target model.
The core of the SMIA technique is a stochastic perturbation method. In each iteration, the attack applies a probabilistic masking operation to low-energy (i.e., quiet) regions of the audio’s frequency spectrum. This strategy ensures the perturbations are perceptually inconspicuous, making the attack virtually undetectable by human listeners. Furthermore, because these modifications are subtle and distributed across the spectrum rather than being concentrated as predictable noise, they are inherently resilient to conventional audio filtering or preprocessing defenses. While each perturbation is minor, they accumulate over iterations to create a significant distortion in the feature space recognized by the models, leading to a successful attack.
Our attack framework is thus composed of two integral components:
  • Iterative Black-Box Optimization: A feedback-driven loop that queries the target model and uses its decision outcome to adaptively update the params to guide the generation of the adversarial audio.
@@
3.2. Iterative Black-Box Optimization
Our attack is operationalized as a black-box iterative search designed to find an adversarial audio sample that simultaneously bypasses both the voice authentication (VAS) and anti-spoofing (CM) systems. The core of this process, detailed in Algorithm 1, employs a multi-modal perturbation strategy. Rather than relying on a single modification technique, the algorithm systematically cycles through a set of distinct perturbation modes: Interpolate, Masking, and a Hybrid approach. For each mode, it iteratively applies subtle modifications to a clone of the original spoofed audio. Based on the feedback from each attempt, the loop updates its core parameters ($T_{dB},\gamma,\sigma,\text{n\_fft},\text{hop\_length}$) to guide the subsequent search. The goal at each step is to produce a variant that the VAS accepts as the target speaker and the CM classifies as bona fide (i.e., not spoofed). This adaptive, multi-pronged search increases the likelihood of discovering a successful adversarial example without requiring any internal knowledge of the target models.
Algorithm 1 Multi-Modal Adaptive Adversarial Attack
1:Enrollment embedding $E$, initial attacker audio $x_{orig}$
2:Max iterations per mode NN
@@ 3.2. Iterative Black-Box Optimization
4:for each modemode in ModesModes do
5:  $x\leftarrow x_{orig}$
6:  Init perturbation params $P\leftarrow\{T_{dB},\gamma,\sigma,\text{n\_fft},\text{hop\_length}\}$
7:  for $i\leftarrow 0$ to N1N-1 do
8:   $x_{adv}\leftarrow\text{ApplyPerturbation}(x,mode,P)$
9:   $vas\_label\leftarrow M_{VAS}(x_{adv},E)$
10:   if $vas\_label=\text{`Accept'}$ then
11:     $cms\_label\leftarrow M_{CMS}(x_{adv})$
12:     if $cms\_label=\text{`Bona Fide'}$ then
13:      return $(x_{adv},\textbf{Success})$
14:     else
15:      $P\leftarrow\text{IncreasePerturbation}(P)$
16:     end if
17:   else
18:     $P\leftarrow\text{DecreasePerturbation}(P)$
19:   end if
20:  end for
21:end for
22:return $(x_{orig},\textbf{Failure: Max iterations reached for all modes})$
@@ 3.3. Spectral Masking and Interpolation
At the core of our attack is a sophisticated spectral perturbation module, ApplyPerturbation. The primary function of this module is to introduce subtle, controlled distortions into an audio signal’s spectrogram. By design, it operates exclusively on low-energy (quiet) time-frequency bins, which are perceptually less significant to the human ear. This strategic targeting is controlled by five key parameters (as shown in Algorithm 2):
  • $T\_{\text{db}}$ A decibel threshold used to identify the quiet parts of the spectrogram.
  • $\gamma$ and $\sigma$ Parameters that control the random selection of bins to be perturbed within the quiet regions.
  • n_fft and hop_length STFT (Short-Time Fourier Transform) settings that define the resolution of the spectrogram grid, which directly influences the interpolation process.
This approach ensures that the resulting adversarial audio remains natural-sounding and free of obvious artifacts. The module employs three primary techniques to manipulate the audio’s underlying acoustic features: masking, interpolation, and a hybrid approach.
Algorithm 2 Spectral Perturbation Module
1:Audio signal $x$, modemode
2:Perturbation params $P=\{T_{dB},\gamma,\sigma,\text{n\_fft},\text{hop\_length}\}$
@@
[…]
9:if $mode=\text{`interpolate'}$ or $mode=\text{`hybrid'}$ then
10:  $\text{mag\_out}\leftarrow\text{Interpolate}(\text{mag\_out},\text{final\_mask})$
11:end if
12:if $mode=\text{`mask'}$ or $mode=\text{`hybrid'}$ then
13:  $\text{mag\_out}[\text{final\_mask}]\leftarrow 0$
14:end if
[…]
16:return $x_{adv}$
@@ 3.3.1. Perturbation Modes
  1. (1)
    It identifies all the reliable, high-energy (non-quiet) points over time, using them as anchors.
  2. (2)
    It constructs a linear interpolation function based on these anchor points.
  3. (3)
    It then uses this function to predict what the values should have been in the targeted quiet locations.
Essentially, it reconstructs the quiet segments based on the audio’s natural temporal contour within that specific frequency. The value $y$ for a target point $x$ between two known anchor points $(x_{0},y_{0})$ and $(x_{1},y_{1})$ is calculated using the standard linear interpolation formula:
@@
This process ensures that the filled-in values are spectrally smooth and plausible, making the alteration difficult to perceive.
@@
3.3.2. Efficacy Against VAS and CMS
This dual-operation approach is highly effective because it simultaneously targets the distinct vulnerabilities of both VAS and CMs. The primary challenge is to bypass the CMS without sacrificing the convincing vocal identity needed to fool the VAS.
  • Deceiving the Voice Authentication System (VAS): The initial spoofed audio, generated by Text-to-Speech (TTS) or Voice Conversion (VC) tools, is already designed to mimic the target speaker’s identity. Our main goal is to preserve this biometric similarity during the attack. By confining perturbations to the low-energy (quiet) regions, we alter the signal’s subtle characteristics, which are scrutinized by a CMS, while leaving the dominant, identity-defining acoustic patterns largely intact. This surgical approach ensures the audio retains the core vocal signature required for the VAS to accept it as the legitimate user.
  • Evading the Countermeasure System (CM): A CM is trained to detect artificial signals by identifying unnatural artifacts, such as vocoder noise or phase discontinuities. Our interpolation technique directly counters this. By filling quiet regions with spectrally consistent data derived from the surrounding audio, we generate a spectrogram that appears smooth and natural. This process removes or masks many of the sharp, anomalous features that a CMS typically flags as evidence of a spoof, allowing the audio to be classified as bona fide.
3.4. Design Justification
  • Quiet Region Targeting: Perturbations in quiet regions are less perceptible to human listeners and allow for stealthy adversarial modifications.
  • Adaptive Strategy: Dynamic adjustment of perturbation parameters improves attack success while maintaining signal intelligibility. Our approach explicitly tests both anti-spoofing and voice authentication systems (VAS). When perturbations are overly strong, similarity with the VAS is lost; when too minimal, the anti-spoofing system cannot be bypassed. This adaptive strategy is well-suited to balance these conflicting objectives.
  • Randomized Selection: Introducing a random element to the selection of bins offers two key advantages. First, it provides greater flexibility regarding the regions of the spectrogram that are modified. Second, by constantly changing the perturbation pattern, it makes the attack significantly harder to detect, as there is no fixed, predictable signature for a defense system to learn.
@@ 4. Experimental Setup
Our experimental setup leverages the LibriSpeech (Panayotov et al., 2015) and ASVSpoof2019 (et al., 2019) datasets to assess system performance. We benchmark our proposed attacks against three state-of-the-art anti-spoofing models: RawNet2 (Tak et al., 2021b), RawGAT-ST(Tak et al., 2021a), and RawPC-Darts(Ge et al., 2021). For the voice authentication stage, we employ three distinct systems: X-Vectors (Snyder et al., 2018), DeepSpeaker (Li et al., 2017), and the commercial Azure Speaker Verification API (Microsoft, [n. d.]). All experiments were conducted on a single NVIDIA T4 GPU. The evaluation is structured to analyze performance under various configurations:

### v3

@@ 2.5.2. Black-box Attacks
In the domain of black-box adversarial attacks, Hai et al. (hai2023sifdetectcracker) proposed SiFDetectCracker, a framework targeting CMs. Their method generates perturbations for speaker-irrelevant features (SiFs), specifically the background noise and the mute parts of an audio sample. The framework was evaluated against several advanced detectors, including RawNet2(rawnet2), RawGAT-ST(rawgat), RawPC-Darts (rawdarts), and Deep4SNet (ballesteros2021deep4snet), using the ASVspoof 2019 dataset (Todisco2019ASVspoof). The attack proved highly effective, achieving an average success rate of 82.2% by showing that many detection systems are sensitive to these non-speech features. While the paper presents a novel attack vector, ablation studies showed that the attack’s performance drops significantly when either time or noise perturbations are removed, indicating a reliance on both factors. Furthermore, the transferability between models with different architectures was low, a limitation our work aims to address.
While the work by Kassis and Hengartner (kassis2023breaking) presents a novel methodology for bypassing VAS, its real-world applicability is limited by key limitations in its evaluation. Their results against combination systems rely on a cumulative metric, which is not descriptive enough to show the state or the actual effectiveness of their attack against these more complex defenses. Furthermore, the attack’s performance dropped significantly when tested against SOTA anti-spoofing systems, like RawPC-Darts (rawdarts) and RawGAT-ST (rawgat), raising doubts about its effectiveness against modern countermeasures. Finally, the designated-app attack vector relies on the significant assumption that an attacker can use a rooted phone to inject audio, a prerequisite that is often blocked by the security measures of financial applications.
@@ 3. Methodology
A combined voice authentication and anti-spoofing system processes a given voice sample $x\in\mathbb{R}^{m}$ to perform two key tasks: SV and CM. The system grants access only when the sample is verified as belonging to the claimed identity and determined to be a genuine, live utterance.
@@
$S(x,e)=\begin{cases}\text{accept}&\text{if }g(x,e)=\text{accept}\land f(x)=\text{real}\\ \text{reject}&\text{otherwise}\end{cases}$
@@
3.1. Threat Model
Our threat model considers an attacker whose primary goal is to gain unauthorized access to a victim’s account on an SV system by mimicking their voice. The success of the attack is defined by a single successful authentication event.
The attacker operates under a black-box assumption, meaning they have no internal knowledge of the system’s architecture, including the specific VAS and CMs employed. The attacker’s capabilities are as follows:
  • Data Collection: They can collect voice samples of the victim from publicly available sources (e.g., social media), requiring a minimum of 10 seconds of audio to initiate the attack.
  • Sample Generation: They can generate a spoofed audio sample using any voice-cloning tool (e.g., Fish Speech APIs (fish-speech-v1.4) or methods used in ASVspoof 2019 (Todisco2019ASVspoof)).
  • Iterative Optimization: The attacker can repeatedly submit generated audio to the authentication endpoint to observe feedback. They use this feedback to iteratively optimize the sample to bypass defenses.
Regarding the feedback loop, we consider two specific system response scenarios:
  • Score and Label: The system returns both a continuous confidence score and the final decision (like Azure SV (microsoft_speaker)).
  • Label Only: The system returns only a binary pass/fail decision.
Our method is effective in both scenarios. The attacker’s actions are considered stealthy, as a successful login does not interfere with or alert the legitimate user. Also, the attack is applicable to both text-dependent and text-independent SV systems.
@@ 3.2. Overview
As shown in Figure 2, the attack pipeline starts by generating a base spoofed audio. This is accomplished by feeding a victim’s voice sample and a target prompt into a voice synthesis model. Our framework is tool-independent—the attacker can employ Fish Speech APIs (fish-speech-v1.4), ASVSpoof 2019 challenge tools (Todisco2019ASVspoof), or any modern voice cloning system.
Second, this synthesized audio serves as the input to our core attack mechanism—the Spectral Masking and Interpolation Attack (SMIA) system. SMIA uses a black-box optimization algorithm to modify the audio using our proposed Spectral Masking and Interpolation technique. The resulting adversarial audio is submitted to the target system (VAS and CM), and the prediction outcome is used as feedback to update the SMIA parameters. This process is repeated, progressively refining the audio until it successfully deceives the target model. Our attack framework is thus composed of two integral components:
  • Black-Box Optimization Framework: A feedback-driven loop that queries the target model and uses its decision outcome to adaptively update the params to guide the generation of the adversarial audio.
@@
3.3. Black-Box Optimization (BBO) Framework
This component serves as the “engine” of the attack, systematically searching for the optimal parameters for the spectral perturbation module (detailed in Section 3.4). As defined in our black-box threat model (Section 3.1), we assume no gradient information or internal model access is available.
To efficiently navigate the search space, we employ a Tree-structured Parzen Estimator (TPE). TPE is a Bayesian optimization technique ideally suited for this task, as it builds probabilistic models of “good” ($l(\theta)$) and “bad” ($g(\theta)$) parameters to suggest candidates that maximize the likelihood of success. Our framework is designed to be flexible, adapting its objective function based on the feedback tuple $(s,l)$ returned directly by the target model, where $s$ is the confidence score and $l$ is the decision label.
Unlike varying threshold-based approaches, our optimization utilizes the target’s output directly. The optimization objective $J(\theta,m)$ is defined adaptively:
@@ 3.3. Black-Box Optimization (BBO) Framework
In the primary scenario where the target returns a continuous score $s$, the TPE optimizer uses this value directly as the scoring function to guide the search toward higher confidence regions. In strict decision-based scenarios, the score collapses to a binary value. Crucially, the returned label $l$ serves as the termination criterion for the attack loop; the optimization halts immediately once a successful label is received, regardless of the score magnitude.
The TPE optimizer models the parameter distributions as shown in Equations 3 and 4, where $J^{*}$ is a quantile of observed objective scores. It suggests new candidates that maximize the ratio $l(\theta)/g(\theta)$, efficiently focusing the search on promising regions.
(3) $l(\theta)$ $=p(\theta|J(\theta,m)>J^{*})$
(4) $g(\theta)$ $=p(\theta|J(\theta,m)\leq J^{*})$
To define the optimizer’s constraints, we bound its search space and budget. The hyperparameter ranges are: the threshold $t_{db}$ is searched between -60.0 dB and -10.0 dB; the probability mean $\mu$ is searched between 0.0 and 1.0; and the probability standard deviation $\sigma_{p}$ is searched between 0.0 and 0.75. The framework performs this for a maximum of $N_{iter}=100$ trials, as detailed in Algorithm 1.
Algorithm 1 Black-Box Optimization Attack
1:Input: Enrolled user ID $U_{id}$, Audio Verification Sample $A_{V}$, Max iterations $N_{iter}$
2:Output: Attack success status and best result $R_{best}$
3:procedure Attack($U_{id}$, $A_{V},N_{iter}$)
4:  $M\leftarrow\{\text{`interpolate', `mask', `hybrid'}\}$ $\triangleright$ Attack modes
5:  $S_{best}\leftarrow$-$\infty$
6:  $R_{best}\leftarrow\text{None}$
7:  $\text{success}\leftarrow\text{False}$
8:  for $m\in M$ do
9:   Initialize BBO sampler $\mathcal{S}_{m}$ $\triangleright$ $\theta=\{t_{db},\mu,\sigma_{p}\}$
10:   for $i\leftarrow 1\text{ to }N_{iter}$ do
11:     $\theta\leftarrow\mathcal{S}_{m}.\text{suggest\_parameters}()$
12:     $A_{C}\leftarrow\textbf{ApplyPerturbations}(A_{V},\theta,m)$
13:     $(s,l)\leftarrow\text{Authenticate}(\textbf{$U_{id}$},A_{C})$
14:     $\mathcal{S}_{m}.\text{report\_result}(s,\theta)$
15:     if $s>S_{best}$ then
16:      $S_{best}\leftarrow s$
17:      $R_{best}\leftarrow(s,l,\theta,m)$
18:     end if
19:     if $l=\text{`success'}$ then
20:      $\text{success}\leftarrow\text{True}$
21:      break
22:     end if
23:   end for
24:   if $\text{success}=\text{True}$ then
25:     break
26:   end if
27:  end for
28:  return $R_{best}$
29:end procedure
@@ 3.4. Spectral Masking and Interpolation
At the core of our attack is a perturbation module, ApplyPerturbation. The function of this module is to introduce subtle, controlled distortions into an audio signal’s spectrogram. By design, it operates exclusively on low-energy (quiet) time-frequency bins, which are perceptually less significant to the ear.
This strategic targeting is controlled by the three key parameters $\theta=(t_{db},\mu,\sigma_{p})$ from our optimizer (as shown in Algorithm LABEL:alg:spectral_perturbation):
  • $t_{\text{db}}$: A decibel threshold used to identify the quiet parts of the spectrogram.
  • $\mu$: The mean probability for selecting a quiet bin.
  • $\sigma_{p}$: The standard deviation of the probability, introducing randomness to the selection.
Our algorithm uses $\mu$ and $\sigma_{p}$ to define the probability $p$ of a quiet bin being perturbed. This value $p$ is sampled from a Normal distribution $N(\mu,\sigma_{p}^{2})$ and then clipped to the valid range $[0,1]$. This stochastic approach (sampling $p$ for each attack) prevents a fixed, detectable perturbation pattern.
This overall method ensures that the resulting adversarial audio remains natural-sounding and free of obvious artifacts. The module employs three primary techniques to manipulate the audio’s underlying acoustic features: masking, interpolation, and a hybrid approach.
Algorithm 2 ApplyPerturbations: Spectral Perturbation Function
1:Audio signal $x$, parameter vector $\theta$, attack mode modemode
2:Adversarial audio $x_{adv}$
@@
4:procedure SampleProbability($\mu,\sigma_{p}$)
5:  $p_{raw}\leftarrow\text{SampleFromNormal}(\mu,\sigma_{p})$
6:  return $\text{Clip}(p_{raw},0.0,1.0)$
7:end procedure
8:procedure ApplyPerturbations($x,\theta,mode$)
9:  $\text{mag},\text{phase}\leftarrow\text{STFT}(x)$ $\triangleright$ STFT: Short-Time Fourier Transform that converts $x$ into time–frequency magnitude and phase
10:  $\text{quiet\_mask}\leftarrow\text{mag}<t_{db}$
11:  $p\leftarrow\text{SampleProbability}(\mu,\sigma_{p})$
12:  $\text{random\_mask}\leftarrow\text{UniformRandom}(\text{shape}(\text{mag}))<p$
13:  $\text{final\_mask}\leftarrow\text{quiet\_mask}\land\text{random\_mask}$
14:  $\text{mag\_out}\leftarrow\text{mag}$
15:  if $mode=\text{`interpolate'}$ or $mode=\text{`hybrid'}$ then
16:   $\text{mag\_out}\leftarrow\text{Interpolate}(\text{mag\_out},\text{final\_mask})$
17:  end if
18:  if $mode=\text{`mask'}$ or $mode=\text{`hybrid'}$ then
19:   $\text{mag\_out}[\text{final\_mask}]\leftarrow 0$
20:  end if
21:  $x_{adv}\leftarrow\text{ISTFT}(\text{mag\_out},\text{phase})$ $\triangleright$ ISTFT reconstructs the time-domain signal from modified spectral components
22:  return $x_{adv}$
23:end procedure
@@ 3.4.1. Perturbation Modes
  • It identifies all reliable, high-energy (non-quiet) points over time, using them as anchors.
  • It constructs a one-dimensional linear interpolation function based on the time indices and magnitude values of these anchor points.
  • It uses this function to generate a new temporal contour for the entire frequency bin, and then selectively replaces the original magnitudes only at the targeted quiet locations with the corresponding values from this new contour.
Essentially, it reconstructs the quiet segments based on the audio’s natural temporal contour within that specific frequency. The magnitude $y$ at a time $x$ falling between two known anchor points $(x_{0},y_{0})$ and $(x_{1},y_{1})$—where $x$ represents time and $y$ represents magnitude—is calculated using the standard linear interpolation formula:
@@
This process ensures that the filled-in values are temporally smooth and plausible, making the alteration difficult to perceive.
@@ 4. Experimental Setup
Our experimental setup leverages the LibriSpeech (panayotov2015librispeech) and ASVSpoof2019 (Todisco2019ASVspoof) datasets to assess system performance. We benchmark our proposed attacks against three state-of-the-art anti-spoofing models: RawNet2 (rawnet2), RawGAT-ST(rawgat), and RawPC-Darts(rawdarts). For the voice authentication stage, we employ three distinct systems: X-Vectors (xvectors), DeepSpeaker (deepspeaker), and the commercial Azure Speaker Verification API (microsoft_speaker). All experiments were conducted on a single NVIDIA T4 GPU. The evaluation is structured to analyze performance under various configurations: